# Robust direction-dependent gain-calibration of beam-modelling errors far from the target field

### v1

@@ 2.2 Direction-dependent gain-calibration with non-smooth beams
When the beam is rapidly varying as a function of frequency, however, the assumptions underlying Equation 4 break down. The most obvious issue is that there may be modelling errors between the true beam and the beam model $\bm{\mathsf{A}}_{p,l,d}$ on scales smaller than a solution interval in frequency and time. In this case, the single $2\times 2$ Jones matrix $\bm{\mathsf{J}}_{p,k}$ computed per interval and direction cannot describe the small-scale beam modelling errors accurately. However, deviations from the beam model within the station typically lead to the primary beam being smoother than the ideal model beam. This is because beam shape variations, and amplitude and phase deviations between elements in a phased station break the station symmetry. Therefore, they tend to blur sharp features in the beam, but also tend to lead to a less sharp fall-off of the amplitude of the far sidelobes, as we illustrate in Sections 3.2 and 4.1. As long as a sufficiently small calibration interval is chosen, the errors introduced by these intra-interval modelling errors are expected to be small, because the beam modelling errors do not lead to more rapidly varying errors than are present in the model beam.
@@
The second, bigger problem, and focus of this work, occurs when the beam varies strongly within a spectral smoothness interval. We illustrate this effect for an extreme case with a single polarisation in Figure 1. In both panels, the right-hand axis (corresponding to the thick lines) shows how the flux of a single point-source that is far from the target direction varies as a function of frequency. The thick tan line corresponds to how the source is ‘seen’ with the true primary beam (with switched-off elements within the station, we refer to this as the ‘true’ beam) and the thick grey line shows the same with an ideal model primary beam (with all elements functioning, we refer to this as the ‘model’ or ‘calibration’ beam). The latter has an area of very high attenuation (a ‘null’), whereas this null is spectrally shifted, wider, and less deep in the true beam.
@@
The left-hand side axes (in regular weight), shows calibration results. In the top panel, the calibrated gains are shown, and in the bottom panel, these gains are multiplied by the source with the calibration beam applied. This is how the beam response is modelled for this station after calibration, and it should ideally coincide with the true response shown by the tan line. In the absence of other errors, the computed gains $\bm{\mathsf{J}}_{p,k}(\nu_{\mathrm{sol}})$ should be similar to the ratio between the true beam and the calibration beam [^8] This statement and figure are only meant to create intuition for the general behaviour of the gain, because the gain computed for each station heavily depends on the modelling errors of the other stations as well.. The ideal gain (that would calibrate this direction perfectly) is shown as the orange line in the top panel.
@@
However, in a noisy observation, the first-iteration least-squares solution per spectral solution interval $\bm{\mathsf{\widetilde{L}}}_{p}(\nu_{\mathrm{sol}})$ looks like the blue line in the top panel. There is a very large spike in the gain at the null in calibration beam. Furthermore, due to the many degrees of freedom, the observational noise is still clearly visible in the least-squares solution. This is normally not a desirable solution, because having too many degrees of freedom leads to overfitting and an undesired interplay between the gains computed for different calibration directions. Additionally, applying noisy solutions to a bright sky model leads to increased errors in the residual data when these sources are subtracted. The product of the gain with the source in the calibration beam (blue line in the bottom panel) therefore does not model the actual source response (thick tan line) well.
@@
When the low-pass filter (i.e. the truncated Gaussian kernel) is applied to constrain the solution, we obtain the green lines. Although the noise on the gain has been reduced, the channels adjacent to the null have a higher gain when regularisation is applied. This does not model the true source well near the null, as is evident from comparing the green and thick tan lines in the bottom panel of Figure 1. In this way, erroneous gains obtained in the null can spread to channels further away from that null. Therefore, although the calibration solutions will generally be improved because of the reduction in degrees of freedom through spectral regularisation, areas with fast spectral beam variation will suffer from gain errors in the channels adjacent to the null.
@@
The first solution that may come to mind is to remove the calibration beam from the algorithm, as this would remove the spike in the ratio between the true and calibration beam. However, this would mean that any beam variations smaller than a calibration interval in time, frequency or space (i.e. a calibration cluster) cannot be modelled. Therefore, a beam model that is correct to the first order is preferred over removing the beam model altogether, although erroneous solutions may need to be excised after calibration. Furthermore, we illustrate the effect here for a large beam error, but the same can occur if the primary beam model is perfect, as long as the null is present. In this case, the spike results from the low signal-to-noise ratio at the null, and the adjacent channels are still perturbed. Excising the parts of the data where these nulls occur is also not a good solution, because it leads to a high loss of data. This is because there can be multiple off-axis sources, so there can be many time and frequency intervals where one of these crosses a null. Furthermore, because this would involve excising entire stations rather than single baselines, many data can be affected.
@@
Figure 1: Illustration of the effect of an erroneous beam model on calibration. In both panels, the bold axis shows the prediction of the response of a station to a single point-source, both with the beam model used in calibration (grey) and the actual perturbed beam (tan). Top: The regular weight axis shows the gain solutions for the same station. Here, blue is the unregularised solution $\bm{\mathsf{J}}_{p,k}(\nu_{\mathrm{sol}})$, green is the same solution after standard smoothing has been applied and red shows the same with the proposed robust regularisation. Orange illustrates the ideal gain that would perfectly match the calibration and true beam. Bottom: the regular weight axis shows the product of the source in the calibration beam and the colored lines from the top panel. For perfect calibration, this aligns with the tan line.
@@ 2.3 Robust spectral regularisation
As an alternative to the current form of spectral regularisation, we propose a modified version of the Gaussian-smoothing algorithm that naturally fits in the current approach through reweighting the regulariser. Effectively, this reweighting scheme changes the spectral smoothness assumption; without reweighting, the assumption is that the gains are smooth in frequency. However, with the reweighted regulariser, we assume that the beam response (i.e. the product of the gain and the calibration beam) is a smooth function of frequency. This is a more accurate assumption, because the true beam has shallower nulls than the calibration beam (e.g. due to element-gain variations and mutual coupling).
@@
for each frequency interval ${\nu_{\mathrm{sol}}}$. The sum in Equation 10 acts as a model-flux weighted estimate of the beam response in direction $k$. The weights are high when the beam-attenuated calibration model has a high absolute value, but low when there are nulls in the calibration beam. Therefore, the high-value gain corrections (the spike in Figure 1) are down-weighted during convolution, such that they have a lower impact on adjacent channels. This leads to a regularized gain solution that is not as sensitive to large deviations. We illustrate the reweighted gain update with the red line in Figure 1. Outside of the channels that fall exactly on the beam null in the calibration beam, the reweighted method (red line) approximates the ideal gain solution (orange line) much better than the unregularised (blue) and standard regularised (green) lines in the top panel. As a result, it outperforms the other methods in approximating the true station response (thick tan line) in the bottom panel.
We refer to the reweighted method as the ‘robust’ weighting scheme (and compare it to the ‘standard’ weighting scheme, which is not reweighted by the model flux) and explain how it is tested on simulated data in the next sections. We do emphasise that this reweighted calibration scheme is robust against errors in the spectral smoothness assumption, but does not remedy errors on scales smaller than a solution interval or calibration cluster. A DD-gain calibration method still divides the sky in calibration clusters, effectively enforcing a top-hat function on the sky, and similarly divides the time axis in discrete solution intervals. The robust calibration scheme is therefore mainly effective against beam modelling errors that change rapidly as a function of frequency.
@@ 3 Simulations
Because beam changes resulting from gain-deviations on the individual antenna element level have not previously been investigated for full-scale LOFAR-HBA observations, we have implemented the simulation and calibration pipeline in a Python code named SHIMMERR (Station Heterogeneity Impact on Multi-dimensional beam-Modelling Errors simulatoR and calibratoR) [^9] https://github.com/Stefanie-B/shimmerr. This codebase is able to simulate any hierarchical interferometer, as long as the European Terrestrial Reference System coordinates of the individual antenna elements are known. It has been highly parallelised to allow for simulations of full LOFAR datasets and can simulate beam errors stemming from broken or miscalibrated (sets of) antenna elements, and pointing errors. When these effects are disabled, the elements at each level of the beam-forming hierarchy become identical, such that their beams only need to be computed once. The computation can then be sped-up by calculating a single beam and simulating the full beam by applying the array factor.
@@ 3.1 Telescope Model
The goal of the forward simulations is to generate a set of ‘true’ visibilities that describe the behaviour of the telescope if there are station-beam perturbations. The basic building block of a LOFAR-HBA station is a cross-bowtie antenna. Sixteen of such antennas form a ‘tile’ (see the first panel in the top row of Figure 3). Tiles are analogue beamformed (van Haarlem et al., 2013). A group of 24, 48, or 96 tiles – depending on the type of station – are digitally beamformed to create a station beam. Station responses to the sky are correlated to create visibilities.
@@
An example of the beams for different beam-forming hierarchy levels within the station are shown in the middle row of Figure 3. In the first column, we show a beam model for the antenna element, based on the beams presented in Chapter 2 of Heald et al. (2018). This beam does not need to be very accurate, because its large width in comparison to the final station-beam width results in limited impact on calibration. The second column shows the array factor of a tile (not multiplied by the element beam). This beam can be steered through analogue beamforming. In this simulation setup, it is steered towards the North Celestial Pole (NCP) field. The third column shows the array factor of the station (with tiles as elements), not multiplied by the tile beams. The final column shows the full station beam. If no perturbations to the individual elements are present, this is equal to the product of the three beams to the left of it.
@@ 3.2 Beam Perturbations
The exact distribution of broken elements is unknown. We simulate beam errors by randomly switching off a number of elements, by setting their element gains in Equation 3 to zero. $N\sim\mathcal{U}(0,3)$ elements are switched off in each tile, where $\mathcal{U}(0,3)$ is the discrete uniform distribution between 0 and 3. This results in an average of 1.5 antennas, or 9 per cent of the tile being simulated as ‘broken’. The effects on the beam, resulting from a random realisation of removing antennas, is illustrated in the last row of Figure 3. As is clear from the second column, switching off three antenna elements drastically alters the shape of the tile beam, removing and altering nulls. Because the tile beams are no longer identical, the nulls and sidelobes of the full station primary beam also change. This has a limited effect near the phase centre (bright yellow spot), but does change when off-axis sources pass through nulls. We illustrate where such off-axis sources can be compared to the phase centre through red circles at the positions of Cas A and Cyg A.
@@ 3.4 Prediction Pipeline
We use the perturbed beam as the propagation matrix in Equation 2 to simulate visibilities. The thermal noise is created by independently drawing the real and imaginary part of the noise from a Gaussian with standard deviation $\sigma=\text{SEFD}/\sqrt{2\Delta\nu\Delta t}$ [Jy] [^11] The factor two is introduced by independently drawing the real and complex part of the noise., where $\Delta\nu$ and $\Delta t$ are the spectral and temporal resolutions of the simulation in Hertz and seconds, respectively. SEFD is the system equivalent flux density in Jansky. For the calibration model, we sum the coherencies per calibration patch to create patch models, and do not add thermal noise.

### v2

@@ 2.2 Direction-dependent gain-calibration with non-smooth beams
When the beam is rapidly varying as a function of frequency, however, the assumptions underlying Equation 10 break down. To explain this, we define the ‘true beam’, i.e. the station’s effective beam with in situ effects (in the case of our simulations, these consist of switched-off antenna elements), and the ‘model beam’ or ‘calibration beam’, which is the expected beam model $\bm{\mathsf{A}}_{p,l,d}$ (without missing elements). The most obvious issue is that there may be modelling errors between the true beam and the beam model on scales smaller than a solution interval in frequency and time. In this case, the single $2\times 2$ Jones matrix $\bm{\mathsf{J}}_{p,k}$ computed per solution interval and direction cannot describe the small-scale beam modelling errors accurately. However, deviations from the beam model within the station typically lead to the primary beam being smoother than the ideal model beam. This is because beam shape variations, and amplitude and phase deviations between elements in a phased station break the station symmetry. Therefore, they tend to blur sharp features in the beam, but also tend to lead to a less sharp fall-off of the amplitude of the far sidelobes, as we illustrate in Sections 3.2 and 4.1. As long as a sufficiently small solution interval is chosen, the errors introduced by these intra-interval modelling errors are expected to be small, because the beam modelling errors do not lead to more rapidly varying errors than are present in the model beam.
@@
The second, bigger problem, and focus of this work, occurs when the beam varies strongly within a spectral smoothness interval. We illustrate this effect for an extreme case with a single polarisation in Figure 1. In both panels, the right-hand axis (corresponding to the thick lines) shows how the apparent flux of a single point-source that is far from the target direction varies as a function of frequency. In this example, we use the position of Cas A from the point-source model discussed in Section 3.3, but use a flat-spectrum source with an intrinsic flux of 2.1 kJy. Cas A has a separation from the target field of $\sim 30^{\circ}$. The primary beam width of a LOFAR-HBA station is $\sim 4^{\circ}$ in the simulated frequency range, placing the source well into the sidelobes of the station beam. The thick tan-coloured line corresponds to how the source is ‘seen’ through the true primary beam of a single station (this is the beam in which elements have been switched off). We denote this as ${{A}^{\prime}}_{\!\!p,l,\mathrm{s}}{I}_{\mathrm{s}}$, where ${{A}^{\prime}}_{\!\!p,l,\mathrm{s}}$ is the true primary beam in the direction of Cas A for a single polarisation, and ${I}_{\mathrm{s}}$ is the intrinsic flux of the source. The thick grey line shows the same with an ideal model primary beam (with all elements functioning), which we denote as ${A}_{p,l,\mathrm{s}}{I}_{\mathrm{s}}$ (without a prime). The latter has an area of very high attenuation (a ‘null’), that is spectrally shifted, wider, and less deep in the true beam ${{A}^{\prime}}_{\!\!p,l,\mathrm{s}}$. We simulate with a monochromatic antenna beam, such that all spectral effects here stem from beamforming, i.e. Equation 8.
@@
The left-hand side axes (in regular weight), show mock calibration results after a single iteration. In the top panel, we illustrate gains ${J}_{p,\mathrm{s}}$ [^9] Because the calibration direction contains only a point source, we use the same subscript for $k$ and $d$. that would result from calibration with different types of regularisation. In the bottom panel, these gains are applied to the source’s model apparent flux with the calibration beam, i.e. ${J}_{p,\mathrm{s}}{A}_{p,l,\mathrm{s}}{I}_{\mathrm{s}}$. Applying the gains for both stations in a baseline creates model visibilities that are corrupted in the same way as the data. These corrupted model visibilities are subtracted from the data to remove the source. Therefore, the calibrated model flux ${J}_{p,\mathrm{s}}{A}_{p,l,\mathrm{s}}{I}_{\mathrm{s}}$ shown in regular weight in the bottom panel should ideally coincide with the true response ${{A}^{\prime}}_{\!\!p,l,\mathrm{s}}{I}_{\mathrm{s}}$ shown by the thick tan-coloured line. With our assumptions, the ideal gain (that would calibrate this direction perfectly) is equal to the ratio between the true beam and the calibration beam ${{A}^{\prime}}_{\!\!p,l,\mathrm{s}}/{{A}^{\prime}}_{\!\!p,l,\mathrm{s}}$. This gain is shown as the orange line in the top panel. Because the true beam ${{A}^{\prime}}_{\!\!p,l,\mathrm{s}}$ is unknown, calibration must estimate this ratio from the noisy observations.
@@
In a noisy observation, the first-iteration least-squares solution per spectral solution interval ${\widetilde{L}}_{p}$ looks like the blue line in the top panel. When no regularisation is applied, this will be the gain after the first iteration ${J}_{p,\mathrm{s}}$. There is a very large spike in the gain at the null in calibration beam. Furthermore, due to the many degrees of freedom, the observational noise is still clearly visible in the least-squares solution, as relatively large spectral channel-to-channel variations throughout the band in the blue line in the top panel. This is normally not a desirable solution, because having too many degrees of freedom leads to overfitting and an undesired interplay between the gains computed for different calibration directions. Additionally, applying noisy solutions to a bright sky model leads to increased errors in the residual data when these sources are subtracted. For unregularised calibration, the value of ${J}_{p,\mathrm{s}}{A}_{p,l,\mathrm{s}}{I}_{\mathrm{s}}$ (blue line in the bottom panel) therefore does not model the actual source response ${{A}^{\prime}}_{\!\!p,l,\mathrm{s}}{I}_{\mathrm{s}}$ (thick tan-coloured line) well.
@@
When the low-pass filter (i.e. the truncated Gaussian kernel) is applied to constrain the solution, we obtain the green lines. Although the noise on the gain has been reduced, the channels adjacent to the null have a higher gain when regularisation is applied. This does not model the true source well near the null, as is evident from comparing the green and thick tan-coloured lines in the bottom panel of Figure 1. In this way, erroneous gains obtained in the null can spread to channels further away from that null. Therefore, although the calibration solutions will generally be improved because of the reduction in degrees of freedom through spectral regularisation, areas with fast spectral beam variation will suffer from gain errors in the channels adjacent to the null.
@@
The first solution that may come to mind is to remove the calibration beam from the algorithm, as this would remove the spike in the ratio between the true and calibration beam. However, this would mean that any beam variations smaller than a solution interval in time, frequency or space (i.e. a calibration cluster) cannot be modelled. Therefore, a beam model that is correct to the first order is preferred over removing the beam model altogether, although erroneous solutions may need to be excised after calibration. Furthermore, we illustrate the effect here for a large beam error, but the same can occur if the primary beam model is perfect, as long as the null is present. In this case, the spike results from the low signal-to-noise ratio at the null, and the adjacent channels are still perturbed. Excising the parts of the data where these nulls occur is also not a good solution, because it leads to a high loss of data. This is because there can be multiple off-axis sources, so there can be many time and frequency intervals where one of these crosses a null. Furthermore, because this would involve excising entire stations rather than single baselines, many data can be affected.
@@
Figure 1: Illustration of the effect of an erroneous beam model on calibration. In both panels, the bold axis shows the prediction of the response of a station to a single point-source, both with the beam model used in calibration (grey) and the actual perturbed beam (tan). Top: The regular weight axis shows the gain solutions for the same station. Here, blue is the unregularised solution ${J}_{p,k}(\nu_{\mathrm{sol}})$, green is the same solution after standard smoothing has been applied and red shows the same with the proposed robust regularisation. Orange illustrates the ideal gain that would perfectly match the calibration and true beam. Bottom: the regular weight axis shows the product of the source in the calibration beam and the colored lines from the top panel. For perfect calibration, this aligns with the tan line.
@@ 2.3 Robust spectral regularisation
As an alternative to the current form of spectral regularisation, we propose a modified version of the Gaussian-smoothing algorithm that naturally fits in the current approach through reweighting the regulariser. Effectively, this reweighting scheme changes the spectral smoothness assumption; without reweighting, the assumption is that the gains are smooth in frequency. However, with the reweighted regulariser, we assume that the beam response (i.e. the product of the gain and the calibration beam) is a smooth function of frequency. This is a more accurate assumption, because the true beam has shallower nulls than the calibration beam, due to variations within the station that break its symmetry. In our simulations, these variations consist of switched-off elements. In general, intra-station variations occur due to a multitude of additional reasons, such as element-gain variations or variations in mutual coupling between antennas at the centre or the edge of a station (Elder & Jacobs, 2024).
@@
for each frequency interval ${\nu_{\mathrm{sol}}}$. The sum in Equation 16 acts as a model-flux weighted estimate of the beam response in direction $k$. The weights are high when the beam-attenuated calibration model has a high absolute value, but low when there are nulls in the calibration beam. Therefore, the high-value gain corrections (the spike in Figure 1) are down-weighted during convolution, such that they have a lower impact on adjacent channels. This leads to a regularized gain solution that is not as sensitive to large deviations. We illustrate the reweighted gain update with the red line in Figure 1. Outside of the channels that fall exactly on the beam null in the calibration beam, the reweighted method (red line) approximates the ideal gain solution (orange line) much better than the unregularised (blue) and standard regularised (green) lines in the top panel. As a result, it outperforms the other methods in approximating the true station response (thick tan-coloured line) in the bottom panel.
We refer to the reweighted method as the ‘robust’ weighting scheme (and compare it to the ‘standard’ weighting scheme, which is not reweighted by the model flux) and explain how it is tested on simulated data in the next sections. We do emphasise that this reweighted calibration scheme is robust against errors in the spectral smoothness assumption, but does not remedy errors on scales smaller than a solution interval or calibration cluster. A DD-gain calibration method still divides the sky in calibration clusters, effectively enforcing a two-dimensional step-function on the sky, and similarly divides the time axis in discrete solution intervals. The robust calibration scheme is therefore mainly effective against beam modelling errors that change rapidly as a function of frequency.
@@ 3 Simulations
Because beam changes resulting from gain-deviations on the individual antenna element level have not previously been investigated for full-scale LOFAR-HBA observations, we have implemented the simulation and calibration pipeline in a Python code named SHIMMERR (Station Heterogeneity Impact on Multi-dimensional beam-Modelling Errors simulatoR and calibratoR) [^10] https://github.com/Stefanie-B/shimmerr. This codebase is able to simulate any hierarchical interferometer, as long as the European Terrestrial Reference System coordinates of the individual antenna elements are known. It has been highly parallelised to allow for simulations of full LOFAR datasets and can simulate beam errors stemming from broken or miscalibrated (sets of) antenna elements, and pointing errors. When these effects are disabled, the elements at each level of the beam-forming hierarchy become identical, such that their beams only need to be computed once. The computation can then be sped-up by calculating a single beam and simulating the full beam by applying the array factor (Equation 8, where $m$ can iterate over antennas or tiles, based on the level in the hierarchy).
@@ 3.1 Telescope Model
The goal of the forward simulations is to generate a set of ‘true’ visibilities that describe the behaviour of the telescope if there are station-beam perturbations. The basic building block of a LOFAR-HBA station is a cross-bowtie antenna. Sixteen of such antennas form a ‘tile’ (see the first panel in the top row of Figure 3). Tiles are analogue beamformed (van Haarlem et al., 2013). A group of 24, 48, or 96 tiles – depending on the type of station – are digitally beamformed to create a station beam. The digitally beamformed station responses to the sky are correlated to create visibilities, as described in Equation 4.
@@
An example of the beams for different beam-forming hierarchy levels within the station are shown in the middle row of Figure 3. In the first column, we show a beam model for the antenna element, based on the beams presented in Chapter 2 of Heald et al. (2018). This beam does not need to be very accurate, because its large width in comparison to the final station-beam width results in limited impact on calibration. The second column shows the array factor of a tile (not multiplied by the element beam). This beam can be steered through analogue beamforming. In this simulation setup, it is steered towards the North Celestial Pole (NCP) field. The third column shows the array factor of the station (with tiles as elements), not multiplied by the tile beams. The final column shows the full station beam. If no perturbations to the individual elements are present, this is equal to the product of the three beams to the left of it. LOFAR-HBA stations are rotated with respect to one another to mitigate the effect of strong sidelobes and nulls (Wijnholds et al., 2011), resulting in differing beam patterns between stations.
@@ 3.2 Beam Perturbations
The exact distribution of broken elements is unknown. Therefore, we simulate beam errors by randomly switching off a number of elements, by setting their element gains in Equation 7 to zero. Throughout our simulations, $N\sim\mathcal{U}(0,3)$ elements are switched off in each tile, where $\mathcal{U}(0,3)$ is the discrete uniform distribution between 0 and 3. This results in an average of 1.5 antennas, or 9 per cent of the tiles being simulated as ‘broken’. We randomise the broken antennas for each tile in each station, such that different stations have a different configuration of broken elements. Because the number of broken antennas per tile is randomised, different stations also have varying numbers of broken elements, as shown in Figure 4.
The effects on the beam, resulting from a random realisation of removing antennas, is illustrated in the last row of Figure 3. As is clear from the second column, switching off three antenna elements drastically alters the shape of the tile beam, removing and altering nulls. Because the tile beams are no longer identical, the nulls and sidelobes of the full station primary beam also change. This has a limited effect near the phase centre (bright yellow spot), but does change when off-axis sources pass through nulls. We illustrate where such off-axis sources can be compared to the phase centre through red circles at the positions of Cas A and Cyg A.
@@
Figure 4: Layout of the stations included in the simulation. The y-axis shows the North-South offset from the centre of the interferometer, while the x-axis shows the East-West offset. The stations are colour-coded by the number of broken antenna elements in the true beam. The total number of antennas in a station is 384.
@@ 3.4 Prediction Pipeline
We use the true beam (with switched-off elements) as the propagation matrix in Equation 4 to simulate visibilities. The thermal noise is created by independently drawing the real and imaginary part of the noise from a Gaussian with standard deviation $\sigma=\text{SEFD}/\sqrt{2\Delta\nu\Delta t}$ [Jy] [^12] The factor two is introduced by independently drawing the real and complex part of the noise., where $\Delta\nu$ and $\Delta t$ are the spectral and temporal resolutions of the simulation in Hertz and seconds, respectively. SEFD is the system equivalent flux density in Jansky. The visibilities created in this way (with the true beam and noise) are treated as the observation’s raw uncalibrated data.
For the calibration model, we subsequently use the model beam (that does not contain broken elements). We sum the coherencies per calibration patch to create model visibilities for each direction, and do not add thermal noise.